OpenStudy (anonymous):

One number is 52 more than twice another. Their product is 52 more than twice their sum. Find the numbers.

OpenStudy (whpalmer4):

one number: \(x\) another number: \(y\) one number is 52 more than twice another: \(x = 52+2y\) their product is 52 more than twice their sum: \(x*y = 52+2(x+y)\) plug the first equation into the second equation and solve...

OpenStudy (anonymous):

0 & -26 58 &3

OpenStudy (whpalmer4):

well, only one of those pairs of numbers is a solution. Which one?

OpenStudy (anonymous):

No, I have to have both...my online work requires two different sets of answers :/ Making things difficult lol

OpenStudy (whpalmer4):

I'm sure that your online system requires both of the numbers that satisfy the problem. You gave 4 numbers, 2 of which do NOT satisfy the problem. And regardless of what the system is asking for, I'm asking you to identify the only pair of numbers which satisfies this: "One number is 52 more than twice another. Their product is 52 more than twice their sum."